OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)?

OpenStudy (anonymous):

slope equation of a line is y=mx+b, m is yfinal-yinitial/xfinal-xinitial.

OpenStudy (anonymous):

y-yfinal=m(x-xfinal)

OpenStudy (anonymous):

finding b, yintial=mxfinal +b

OpenStudy (anonymous):

4x - y = 1 4x + y = 1 x - 4y = -13 4x - y = -7 which one

OpenStudy (anonymous):

im not getting any of those answers.

OpenStudy (anonymous):

equation of a line, "y=mx+b" you need to find m and b